### 01 Mar 2024 v1

# DAMS-DETR: Dynamic Adaptive Multispectral Detection Transformer with Competitive Query Selection and Adaptive Feature Fusion

Abstract
Infrared-visible object detection aims to achieve robust even full-day object detection by fusing the complementary information of infrared and visible images. However, highly dynamically variable complementary characteristics and commonly existing modality misalignment make the fusion of complementary information difficult. In this paper, we propose a Dynamic Adaptive Multispectral Detection Transformer (DAMS-DETR) based on DETR to simultaneously address these two challenges. Specifically, we propose a Modality Competitive Query Selection strategy to provide useful prior information. This strategy can dynamically select basic salient modality feature representation for each object. To effectively mine the complementary information and adapt to misalignment situations, we propose a Multispectral Deformable Cross-attention module to adaptively sample and aggregate multi-semantic level features of infrared and visible images for each object. In addition, we further adopt the cascade structure of DETR to better mine complementary information. Experiments on four public datasets of different scenes demonstrate significant improvements compared to other state-of-the-art methods. The code will be released at https://github.com/gjj45/DAMS-DETR.
Keywords:
Object detection Multispectral detection Infrared DETR Query selection Adaptive feature fusion
1 Introduction
Object detection is a fundamental task in computer vision and most of research works are based on visible images with detailed object information, e.g., texture and color information. Thanks to the development of deep learning, the object detection technique has made great progress. However, it is still challenged by poor imaging conditions, such as low illumination, smoke, fog, and so on, which could make objects pretty obscure and further obviously degrade the performance of object detection.
Figure 1: Illustrations of two typical challenges in infrared-visible object detection. (a) Three pedestrians represent different complex complementary characteristics. In this example, the objects in the visible image provide unuseful interference information (red), partial complementary information (blue), and full complementary information (green). (b) One example of the misalignment problem, in which the ground truths of infrared and visible objects appear obvious dislocation. This misalignment commonly happens in infrared-visible images. We propose a Multispectral Transformer Decoder with Multispectral Deformable Cross-attention module to simultaneously address these two typical challenges.
Thus, infrared images are introduced into the object detection task. Different from visible imaging, infrared imaging captures the thermal radiation of objects, making it unaffected by illumination, smoke and fog occlusion conditions. Therefore, infrared imaging can still well capture objects even in low illumination, heavy smoke, or fog, whereas the detailed texture and color information will be lost. These complementary characteristics of infrared and visible imaging not only can improve the performance of object detection, but also is considered to be promising to implement full-day object detection. Thus, infrared-visible object detection has attracted extensive attention in recent years [41, 39, 5, 30, 15, 33].
However, these methods tend to neglect the modality interference encountered in complex scenes during the fusion process. For the case that the object signal in one modality is poor or absent, directly fusing the information of two modalities will bring in unuseful interference information, which could lead to feature confusion and thus degrade the object detection performance. For example, as shown in Fig. 1(a), the pedestrian within the smoke fully disappears, and intuitively, the best way is to suppress or discard the visible information for that pedestrian. Some works learn a global fusion weight to adapt to specific scenes, the representative ones of which are to adopt the illumination-aware network to obtain illumination score as the global fusion weight [12, 41, 33]. Other works learn local region fusion weights through bounding-box level semantic segmentation [12, 41, 33], or regions of interest (ROI) prediction [40, 39, 9].
Actually, due to the fully different imaging principles, the complementary characteristics in infrared-visible images appear highly variable with the specific scenes and objects, as shown as Fig. 1. Especially, from Fig. 1(a), we can observe that three pedestrians have obviously different complementary characteristics. The one with the green bounding box has good complementary information in both modalities, while the one with the red bounding box has only infrared information available, as mentioned previously. In contrast, the one with the blue bounding box has partial information available in both modalities, which commonly exists in practical applications. This situation would make current methods fail to effectively fuse features, even for above mentioned region-based weight fusion methods in which the segmented or predicted regions are usually bigger than objects. Therefore, more fine-grained two-modality information fusion still remains a challenge.
Another important challenge in infrared-visible object detection is the modality misalignment problem. Most feature fusion methods assume that the two modalities are well-aligned. However, precise registration is difficult because infrared-visible images often exhibit significant visual differences and are not always captured at the exact same timestamp [43]. As a result, even through manual registration, the imaging objects in two modalities for the same one are usually misaligned, as shown in Fig. 1(b). This could lead to disrupting the consistency of fused feature representation of current methods, affecting the final detection performance. AR-CNN [40, 39] explicitly learned the offsets of objects in both modalities to achieve alignment on object features. However, this method requires additional paired bounding-box annotations of two modalities during training, which is time-consuming and labor-intensive.
In this paper, we propose a novel adaptive infrared-visible object detection method based on DETR [42, 35, 19]. Inspired by the deformable cross attention [42], we propose a Multispectral Transformer Decoder with Multispectral Deformable Cross-attention module to simultaneously address the above two challenges. Specifically, we adopt an effective strategy of adaptive sparse feature sampling and weight aggregation on two-modality feature maps of different semantic levels. This strategy can effectively fuse fine-grained complementary information even when two modalities are misaligned. Since the two challenges of fine-grained information fusion and modality alignment are simultaneously handled in a single module, our method is more efficient than existing methods which usually handle them separately. Furthermore, unlike the one-step fusion strategy adopted by existing methods, the information fusion of each specific object in our method happens on different semantic levels, which makes the complementary information fully be mined and utilized. Actually, we observed that the complementary information of two modalities also dynamically varies with the semantic levels, as discussed in Sec. 3.3. This is similar to our observation on scenes and objects discussed previously. Thus, our adaptive multi-level fusion is more reasonable.
In order to provide reliable input at the early stage, we design a Competitive Query selection strategy to select dominant modality features for each object as initial position and content queries to build a basic salient feature representation for the Multispectral Transformer Decoder, which can provide useful prior information for following processing. To further exploit more reliable and comprehensive complementary information step by step, the cascaded layer structure of DETR is employed in this paper. Totally, our method is similar to the human observation pattern which dynamically focuses on objects in each modality and gradually aggregates key information of two modalities.
Our contributions can be summarized as follows:
  • 1.
    We propose a novel infrared-visible object detection method, named DAMS-DETR, which can dynamically focus on dominant modality objects and adaptively fuse complementary information.
  • 2.
    We propose a Competitive Selection strategy for multimodal initialization queries to dynamically focus on the dominant modality of each object and provide useful prior information for following fusion process.
  • 3.
    We propose a Multispectral Deformable Cross-attention module, which can simultaneously adaptively mine fine-grained partial complementary information at different semantic levels and adapt to modality misalignment situations.
  • 4.
    Experiments on four public datasets with different scenarios demonstrate that the proposed method achieves significant improvement compared with other state-of-the-art methods.
2 Related Work
Infrared-Visible object detection. Previous research in infrared-visible object detection are primarily built upon the single modality object detection frameworks, which are generally divided into one-stage object detectors, such as Faster RCNN [28], and two-stage object detectors, such as YOLO [25, 26, 27, 31].
In order to fuse the complementary information of infrared and visible images, Konig et al. [10] introduced a fully convolutional fusion RPN network, which fused infrared and visible image features by concatenation, and concluded that halfway fusion can obtain better result [16]. On this foundation, some studies designed CNN-based attention modules to better exploit the potential complementation of infrared and visible images [23, 29, 2]. Additionally, other works introduced transformer-based fusion modules to capture a more global complementary relationship between infrared and visible images [22, 5, 43, 30].
In addition to the above methods of directly fusing image features. Some works adopted illumination information as global weights to fuse infrared and visible image features or post-fuse the multibrance detection results to reduce the impact of interfering information [12, 41, 33]. Considering that the complementary characteristics of different regions could be different, some studies introduced bounding box level semantic segmentation [11, 12, 1, 37, 38] or regions of interest (ROI) prediction [40, 39, 9] to guide the fusion of different regions. Other works also utilized the confidence or uncertainty scores of regions to post-fuse the predictions of multibranchs [14, 15].
To address the challenge of modality misalignment, Zhang et al. [40, 39] developed the AR-CNN network and explicitly aligned the features of two modalities by incorporating additional paired bounding box annotations to learn object misalignment. Kim et al. [8] also employed a multi-label learning approach to adapt object detection in scenes with misalignment.
The above methods significantly enhance the performance of infrared-visible object detection. However, these methods perform overall image feature fusion or one-step reweighting region feature fusion, making it difficult to mine complete complementary information in complex scenes. In contrast, our proposed method links the misalignment problem with complementary feature fusion, which gradually and adaptively mines object-specific fine-grained complementary information at multiple semantic levels.
End to end Object Detectors. In recent years, Carion et al. [3] first introduced the end-to-end object detector based on transformer called DEtection TRansformer(DETR). It views object detection as a set prediction problem and uses binary matching to directly predict one-to-one object sets during training. This greatly simplifies the object detection pipeline, eliminating the need for manual anchor box design or post-processing with NMS. Although DETR has these advantages, it has the problem of slow training convergence and many DETR variants have been proposed to address this issue. Deformable DETR [42] accelerated training convergence by predicting 2D anchor points and designing a Deformable cross-attention module to sparsely sample features around reference points. Conditional DETR [20] decoupled the content and position operations and proposes conditional cross-attention to accelerate training convergence. Efficient DETR [34] proposed a more efficient pipeline by combining dense detection and sparse set detection. DAB-DETR [18] introduced 4D reference points to optimize the prediction box layer by layer. DN-DETR [13] accelerated the training process and label-matching effect by introducing query denoising in the training phase. DINO [35] integrated the above works to build a powerful DETR detection framework. For detection efficiency, RT-DETR [19] constructed a real-time and end-to-end object detector by designing an Efficient Hybrid Encoder and adopting an IoU-aware Query Selection strategy.
Recently, a study on multispectral pedestrian detection based on DETR was conducted, which designed three prediction branches and an instance-aware modality-balance loss to align the contributions of each modality [32]. In contrast, our method has only one prediction branch and guides the feature fusion of each specific object through a dynamic Modality Competitive Query Selection strategy. This dynamic guidance strategy based on sample changes could be more effective for infrared-visible object detection in complex and changeable scenes.
3 Method
3.1 Overview
The overview of our DAMS-DETR is illustrated in Fig. 2. Our method contains four main components: two modality-specific CNN backbones, two modality-specific Efficient Hybrid Encoders,
Figure 2: Overview of DAMS-DETR. Our DAMS-DETR comprises four main components: two modality-specific CNN backbones to extract features, two modality-specific Efficient Hybrid Encoders [19] to encode features, a Modality Competitive Query Selection module to select initial object queries, and a Multispectral Transformer Decoder to mine complementary information.
Modality Competitive Query Selection, and a Multispectral Transformer Decoder. Given a pair of infrared and visible images, we first extract and encode their features separately using two modality-specific CNN backbones and two modality-specific Efficient Hybrid Encoders. Subsequently, the encoded features are flattened, concatenated, and input into the Modality Competitive Query Selection module. This module selects salient modality features to serve as the initial object queries. Next, these modality-specific object queries enter the Multispectral Transformer decoder, which refines them with multiple semantic levels of infrared and visible feature maps through cascaded decoder layers. Finally, these refined object queries are mapped through the detection head to obtain the bounding boxes and classification scores of all objects.
The Efficient Hybrid Encoder we adopted is from RT-DETR [19]. This encoder combines Transformer and CNN to significantly reduce computational complexity. In the following, we will elaborate on the proposed Modality Competitive Query Selection strategy and Multispectral Transformer Decoder with the Multispectral deformable cross-attention module in detail.
3.2 Modality Competitive Query Selection
The object queries in DETR are a set of learnable embeddings, which contain the content and position information of the objects. These queries serve as object feature representations that interact with the image feature sequence in the decoder and generate bounding boxes and classification scores through the prediction head mapping. In addition to setting object queries as learnable embeddings, there are also some methods to use the Top-$K$ score features as the initial object query [42, 35, 34]. Learnable object queries are difficult to optimize because they have no explicit physical meaning [19]. In infrared and visible images, a gap exists between two modality features, further complicating the optimization of learnable object queries. Therefore, selecting object queries from the encoded feature maps is more suitable for the dynamic changes of complementary characteristics in the infrared-visible object detection task.
Concretely, we concatenate encoded feature sequences from the infrared and visible modalities and feed them into a linear projection layer to obtain feature point scores. From this combined feature representation,
Figure 3: Visualization of Modality Competitive Query Selection results. Red points indicate high-score queries selected in the corresponding modality image, while blue points represent lower-scoring queries. The red boxes indicate the objects represented by high-score queries.
we select the Top-$K$ scoring features as the initial object queries. These Top-$K$ features are sourced either from the infrared or visible features, respectively, and each feature represents an object instance specific to their respective modality. This approach can be defined as follows:
$z=\operatorname{Top-\mathit{K}}(\operatorname{Linear}(\operatorname{concat}(I,V))),\vspace{-5pt}$ (1)
where $z$ denotes the set of $K$ selected features, $I$ and $V$ represent the flattened encoded infrared and visible feature sequence, respectively.
As mentioned before, infrared or visible images may contain unuseful interfering information, which could confuse the network. We competitively select modality-specific features to build a salient feature representation for each object. This approach helps to prevent the introduction of interference from another modality in the early stages and provides useful prior information for refining the query in the subsequent decoder, emphasizing that the query’s representation of the object should prioritize the modality from which it originates. Additionally, we use the optimization strategy of IOU-aware classification loss [36, 19] to further improve the quality of the selected features.
Effectiveness analysis. To observe the performance of our Modality Competitive Query Selection strategy in the network, we visualize the positions and scores of the selected modality-specific initial queries on paired images. Specifically, we map these selected features to obtain the coordinates of reference points, which are further projected onto either infrared or visible images. As illustrated in Fig. 3, the visualization shows that different object instances are saliently represented by different dominant modality features. This selection result is consistent with our intuition, demonstrating that this approach dynamically chooses the dominant detection modality for each object under varying conditions. More detailed quantitative results and analysis are presented in Sec. 4.4.
Redundant queries. We also observe that there are redundant queries in infrared and visible images to point to the same object. However, the network effectively eliminates this redundant information thanks to the DETR’s one-to-one matching optimization pattern and the self-attention mechanism applied to all modality-specific queries in the decoder. To further address this potential issue, we introduce Noise Query Learning [35, 13] strategy during training to facilitate learning the optimal modality match for each object.
Figure 4: The structure of the Multispectral Transformer Decoder (DeNoising Training Group is omitted in the figure) and Multispectral Deformable cross-attention module.
3.3 Multispectral Transformer Decoder
The details of the Multispectral Transformer Decoder are illustrated in Fig. 4. In each layer, modality-specific object queries first undergo multi-head self-attention to obtain contextual information and reduce redundancy. Subsequently, our Multispectral Deformable Cross-attention module refines these modality-specific queries with multi-semantic infrared and visible features. Additionally, we employ 4D anchor boxes to constrain the sampling range within the Multispectral Deformable Cross-attention module, and iteratively refine queries and anchor boxes through cascaded decoder layers. Concretely, in the Multispectral Decoder with $D$ layers, we map from the $q$-th modality-specific query $z_{q}^{d}$ in the $d$-th layer to obtain the refined 4D reference point $b_{q}^{d}$. The process can be described as follows:
$b_{q\{x,y,w,h\}}^{d}=\sigma\left(MLP^{d}\left(z_{q}^{d}\right)+\sigma^{-1}\left(b_{q}^{d-1}\right)\right),$ (2)
where $d\in\{2,3,\ldots,D\}$, $MLP$ consists of two linear projection layers, $\sigma$ represents sigmoid function, $\sigma^{-1}$ represents inverse sigmoid function, and $b_{q}^{1}$ is the initialized anchor boxes. The initial anchor setting is consistent with the two-stage method in Deformable DETR [42]. This refined 4D reference point serves as the reference position constraint for subsequent sampling on multi-semantic infrared and visible feature maps in the Multispectral Deformable Cross-attention module.
Multispectral Deformable Cross-attention module. In Deformable DETR [42], key features are aggregated by sparse sampling on the feature map, and we extend it to a multi-modal form to achieve adaptive infrared and visible feature fusion. The detailed architecture of the Multispectral Deformable Cross-Attention module is illustrated in Fig. 4. Specifically, we map the 4D reference points to position embeddings through an MLP layer. After combining the modality-specific query feature with position embeddings, two linear layers are employed to predict the sampling offsets and aggregation weights on the multi-semantic feature maps in both modalities, respectively.
Figure 5: Visualization of feature sampling at different semantic levels in different decoder layers. We use different colors of points to represent the results of sampling in different semantic layers, where blue, green, and red represent sampling points on low-level, middle-level, and high-level semantic features maps respectively. Brightly colored and large points indicate relatively high attention weights. (a) The green boxes in the visible image represent aligned bounding boxes, which show the sampling points in each modality are concentrated on the right instance locations. (b) Objects occluded by background and smoke tend to predominantly focus on the infrared modality at each decoder layer, whereas undisturbed objects simultaneously attend to complementary information from both modalities.
Finally, these sampled multi-semantic infrared and visible features are aggregated by aggregation weights. Due to the ability of this method to predict sampling position offsets independently in both the infrared and visible modalities, the network could still focus on the features of misaligned objects in misaligned image pairs.
Given the input multi-semantic infrared and visible feature maps $\left\{x_{1}^{l},x_{2}^{l}\right\}_{l=1}^{L}$, we use the normalized centerpoint of $b_{q}$ as the 2D reference point $\hat{p}_{q}$. We define the Multispectral Deformable Cross-attention module $F$ as follows:
$\begin{split}&\mathit{F}\left(z_{q},\hat{p}_{q},\left\{x_{1}^{l},x_{2}^{l}\right\}_{l=1}^{L}\right)=\sum_{h}^{H}W_{h}\\[-8.0pt] &\left[\sum_{m}\sum_{l=1}^{L}\sum_{k=1}^{K}A_{mhlqk}\cdot W_{h}^{\prime}x_{m}^{l}\left(\Phi_{l}\left(\hat{p}_{q}\right)+\Psi_{l}\left(\Delta\mathrm{p}_{mhlqk}\right)\right)\right],\end{split}$ (3)
where $m\in\{1,2\}$ denotes the visible and infrared modalities, $h$ indexes the attention head, $l$ indexes the input feature semantic level, and $k$ indexes the sampling point. $A_{mhlqk}$ and $\Delta\mathrm{p}_{mhlqk}$ denote the $k$-th attention weight and sampling point in the $l$-th feature semantic level and the $h$-th attention head within the $m$ modality, respectively. The attention weight $A_{mhlqk}$ is normalized by $\sum_{m}\sum_{l=1}^{L}\sum_{k=1}^{K}A_{mhlqk}=1$. Function $\Phi_{l}\left(\hat{p}_{q}\right)$ scales $\hat{p}_{q}$ to the $l$-th semantic level feature map and function $\Psi_{l}\left(\Delta\mathrm{p}_{mhlqk}\right)$ constrains the predicted offset within the range of $b_{q}$, so as to focus on the information around the object and reduce the difficulty of optimization.
Effectiveness analysis. To observe the effectiveness of this feature fusion approach, we visualize the positions and weights sampled in both modalities at different semantic levels in different decoder layers, as shown in Fig. 5. It shows that as the decoder layer deepens, our method tends to adaptively focus on low-level semantic features in infrared modality and additional high-level semantic features in visible modality. This result is reasonable since the infrared modality carries less information and could provide reliable low-level semantic information such as basic contours and shapes, while the visible modality with more information could additionally provide more abstract high-level semantic information, such as more reliable contextual relationships of object categories. Additionally, We observe that these points can adapt to misalignment scenes and adaptively focus on the key information of the object, such as the edge information, which is important to define the boundary of the object. The differences in sampling positions and weight distribution in different decoder layers also verify the effectiveness of the cascade structure for reliable complementary information mining. More details can be found in Sec. 4.4.
3.4 Loss Function
The training loss of our model follows the DETR-like detectors, which is defined as:
$\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{box}+\mathcal{L}_{dn},\vspace{-2pt}$ (4)
where $\mathcal{L}_{cls}$ is the IoU-aware classification loss following RT-DETR [19], $\mathcal{L}_{box}$ is composed of L1 loss and generalized IoU loss for bounding box regression and $\mathcal{L}_{dn}$ is the loss for denoising training [35]. In addition, we also calculate the loss of each decoder layer as the auxiliary optimization loss.
4 Experiments
4.1 Dataset and metric
[…]
FLIR. We use the aligned version [37] with a resolution of 640 × 512, which contains 4,129 pairs for training and 1,013 pairs for testing and has three object categories. This dataset contains both day and night scenes, with obvious misalignment of image pairs.
LLVIP. The dataset is for pedestrian detection in low-light surveillance scenes. It consists of 12,025 image pairs for training and 3,463 pairs for testing with a resolution of 1280 × 1024, with good image pair registration. Most of the scenes are in dark condition, with only one category of pedestrians.
[…]
Model Backbone Data Type $m$AP50 $m$AP75 $m$AP Inference Speed(s)
Yolov7 CSPDarknet53 IR 60.5 - 34.8 0.014
Yolov7 CSPDarknet53 RGB 69.0 - 42.7 0.014
[…]
DINO ResNet50 IR 58.8 36.1 35.0 0.155
DINO ResNet50 RGB 73.3 48.2 46.3 0.155
[…]
CFT [22] CSPDarknet53 IR+RGB 68.2 44.6 42.5 0.050
ICAFusion [30] CSPDarknet53 IR+RGB 67.8 44.5 41.9 0.033
Ours ResNet50 IR+RGB 80.2 56.0 52.9 0.117
Table 2: Comparisons on the FLIR-aligned Dataset.
Model Backbone Data Type $m$AP50 $m$AP75 $m$AP
Yolov5 CSPDarknet53 IR 80.1 - 42.4
Yolov5 [22] CSPDarknet53 RGB 67.8 25.9 31.8
DINO ResNet50 IR 80.6 42.7 44.8
DINO ResNet50 RGB 70.9 25.9
GAFF [38] ResNet18 IR+RGB 72.9 32.9 36.6
SMPD [15] VGG16 IR+RGB 73.6 - -
CFT [22] CSPDarknet53 IR+RGB 78.7 35.5 40.2
ICAFusion [30] CSPDarknet53 IR+RGB 79.2 36.9 41.4
MFPT [43] ResNet50 IR+RGB 80.0 - -
LRAF-Net [5] CSPDarknet53 IR+RGB 80.5 - 42.8
Ours ResNet50 IR+RGB 86.6 48.1 49.3
VEDAI. The dataset is the multispectral aerial imagery dataset for vehicle detection, consisting of 1,200 image pairs with a resolution of 1024 × 1024, and its image pairs are strictly registered. It includes nine object categories and most of the objects are small, which poses a great challenge for DETR-like detectors. We convert its bounding boxes to horizontal boxes, following [23].
4.2 Implementation Details
We employ ResNet50 [6] as the backbone for both the infrared and visible branches, with a feature map semantic level of L = 3. The Efficient Hybrid Encoder contains one layer, while the MS-Decoder contains six layers. We set the number of attention heads, sampling points, and selected queries as H = 8, K = 4, and N = 300, respectively.
[…]
4.3 Comparison with SOTA
[…]
Table 3: Comparisons on the LLVIP Dataset.
Model Backbone Data Type $m$AP50 $m$AP75 $m$AP
Yolov5 [7] CSPDarknet53 IR 96.5 76.4 67.0
Yolov5 [7] CSPDarknet53 RGB 90.8 56.4 52.7
DINO ResNet50 IR 96.6 70.6 62.9
DINO ResNet50 RGB 91.6 58.0 53.8
CSAA [2] ResNet50 IR+RGB 94.3 66.6 59.2
CFT [22] CSPDarknet53 IR+RGB 97.5 72.9 63.6
LRAF-Net [5] CSPDarknet53 IR+RGB 97.9 - 66.3
MS-DETR [32] ResNet50 IR+RGB 97.9 76.3 66.1
Ours ResNet50 IR+RGB 97.9 79.1 69.6
Table 4: Comparisons on the VEDAI Dataset.
Model Backbone Data Type $m$AP50 $m$AP
Yolo-Fine [21] CSPDarknet53 IR 75.2 -
Yolo-Fine [21] CSPDarknet53 RGB 76.0 -
DINO ResNet50 IR 78.4 41.7
DINO ResNet50 RGB 81.3 44.9
Yolo Fusion [23] CSPDarknet53 IR+RGB 78.6 49.1
CFT [22] CSPDarknet53 IR+RGB 85.3 56.0
LRAF-Net [5] CSPDarknet53 IR+RGB 85.9 59.1
Ours ResNet50 IR+RGB 91.5 55.3
[…]
Comparisons on FLIR: As shown in Tab. 2, our method outperforms single-modality methods and surpasses the best feature fusion methods by 6.1% on $m$AP50 and 6.5% on overall $m$AP, demonstrating that our method can effectively mine complementary information under different illumination conditions.
Comparisons on LLVIP: As shown in Tab. 3, the dataset has nearly saturated performance on $m$AP50, but our method still outperforms the best of $m$AP75 (MS-DETR [32]) by 2.8%, and the best of $m$AP (LRAF-Net [5]) by 3.3%, demonstrating our method could mine more comprehensive and fine-grained complementary information.
Comparisons on LLVIP: As shown in Tab. 3, the dataset has nearly saturated performance on $m$AP50, but our method still outperforms the best of $m$AP75 (MS-DETR [32]) by 2.8%, and the best of $m$AP (LRAF-Net [5]) by 3.3%, demonstrating our method could mine more comprehensive and fine-grained complementary information.
[…]
Effect of Modality Competitive Query Selection (MCQS). We apply standard query selection on the added encoded features to show the effectiveness of Modality Competitive Query Selection. As shown in the 5th and 6th rows of the Tab. 5, our Modality Competitive Query Selection strategy bring 1.1% $m$AP50 and 0.7% $m$AP improvement due to avoiding the early introduction of interference by dynamically selecting modality-specific queries.
Effect of Multispectral deformable cross-attention module (MDCA). We apply standard deformable cross-attention to the fused features obtained by adding the outputs of two encoders in the decoder.
[…]
Modality Module & Strategy Metrics
Vis Ir MCQS MDCA CQS $m$AP50 $m$AP75 $m$AP
1 $\checkmark$ $\checkmark$ 75.7 51.9 49.3
2 $\checkmark$ 75.3 52.0 49.4
3 $\checkmark$ $\checkmark$ 62.2 40.1 38.5
4 $\checkmark$ 62.5 39.9 39.0
5 $\checkmark$ $\checkmark$ $\checkmark$ 77.8 56.0 51.6
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 78.9 55.5 52.3
7 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 79.4 55.8 52.5
8 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 79.3 55.8 52.6
9 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 80.2 56.0 52.9
As shown in the 5th and 7th rows of the Tab. 5, our method brings improvements of 1.5% $m$AP50 and 0.9% $m$AP. This improvement can be attributed to our method’s ability of effectively adapting to modality misalignment scenes while performing adaptive feature aggregation for each object in a more fine-grained way. Certainly, when using both Modality Competitive Query Selection and the Multispectral Deformable Cross-Attention module, we achieved the best results on all AP metrics, as shown in the 9th row of the Tab. 5.
Effect of Content Query Selection (CQS). We analyze the performance of different content query strategies in infrared-visible object detection. For comparison, we take the single branch of our network as the single-modality detection method. In the 8th and 9th rows of the Tab. 5, without CQS means setting the content query as a learnable query, while the position query is still obtained by Modality Competitive Query Selection (similar to mixed query selection in DINO [35]). We can see that using CQS leads to better results, as it provides stronger prior information for subsequent multispectral cross-attention in the Multispectral Decoder. However, as shown in the first 4 rows of Tab. 5, the selection or learnable of content query has a limited impact on single-modality object detection, since the feature representation of single modality is relatively consistent.
4.5 Limitation
Our method is built upon DETR, so the accurate detection performance on small objects may not be as competitive as other CNN-based detectors, as shown in Tab. 4. This is attributed to the fact that the transformer model prioritizes global information, but it is less effective in extracting local information compared to the CNN model, resulting in less precise boundaries for small object bounding boxes. Additionally, our method is effective in addressing common misalignment in the majority of cases. However, it may not handle extreme misalignment situations well, as the object features of the modality will be lost when objects exceed the range of the 4D reference point.
5 Conclusion
In this paper, we propose DAMS-DETR to simultaneously address complementary information fusion and modality misalignment problems in infrared-visible object detection. Through Modality Competitive Query Selection, DAMS-DETR can dynamically choose salient modality feature representations for specific objects based on complementary characteristics. In the Multispectral Deformable Cross-attention module, we link feature fusion and modality misalignment to mine reliable complementary information at multi-semantic levels. Experiments on four datasets with different scenarios demonstrate that the proposed method achieves significant improvements compared to other state-of-the-art methods. We hope this work can inspire more DETR-based infrared-visible object detection work.